### v1

# Inspection-Guided Randomization: A Flexible and Transparent Restricted Randomization Framework for Better Experimental Design

Abstract
Randomized experiments are considered the gold standard for estimating causal effects. However, out of the set of possible randomized assignments, some may be likely to produce poor effect estimates and misleading conclusions. Restricted randomization is an experimental design strategy that filters out undesirable treatment assignments, but its application has primarily been limited to ensuring covariate balance in two-arm studies where the target estimand is the average treatment effect. Other experimental settings with different design desiderata and target effect estimands could also stand to benefit from a restricted randomization approach. We introduce Inspection-Guided Randomization (IGR), a transparent and flexible framework for restricted randomization that filters out undesirable treatment assignments by inspecting assignments against analyst-specified, domain-informed design desiderata. In IGR, the acceptable treatment assignments are locked in ex ante and pre-registered in the trial protocol, thus safeguarding against p-hacking and promoting reproducibility. Through illustrative simulation studies motivated by education and behavioral health interventions, we demonstrate how IGR can be used to improve effect estimates compared to benchmark designs in group formation experiments and experiments with interference.
1 Introduction
Randomized experiments are considered the gold standard for estimating causal effects of interventions. By randomizing the treatment, the distribution of baseline covariates is similar in expectation across treatment arms, and observed differences in the outcome can be attributed to the intervention. However, randomization alone does not guarantee covariate balance for every realized allocation of treatments (Fisher,, 1935; Rubin,, 2008). Suppose the treatment allocation chosen for an experiment, by chance, exhibits imbalance on covariates. At best, the analyst makes careful ex post adjustments for imbalance, usually according to a regression model that relies on assumptions that may or may not hold (Freedman,, 2008). At worst, covariate imbalance is overlooked or handled improperly (e.g., in an attempt to p-hack), leading to unreliable conclusions about treatment effects (e.g., false positives). Restricted randomization helps prevent imbalanced treatment allocations. In restricted randomization, the assignment mechanism is modified such that the probability of choosing an undesirable allocation is set to zero. While performing restricted randomization is rather common, it is primarily applied to the canonical use case of balancing two-arm trials where the target effect estimand is the average treatment effect.
Experimental settings with different design desiderata and estimands could also stand to benefit from a restricted randomization design. A natural extension of balancing a two-arm trial is to balance a multi-arm trial, where the goal is to simultaneously estimate treatment effects for multiple interventions. Additionally, in experiments where we want to explore treatment effect moderators (VanderWeele,, 2015), we may want to construct treatment arms with intentional “imbalance” on a hypothesized moderating covariate but maintain balance on other covariates. We may also be interested in controlling factors beyond balance that could jeopardize the credibility of effect estimates. Causal inference in the presence of interference (Cox,, 1958; Rosenbaum,, 2007), when one individual’s treatment affects a different individual’s outcome, has been an area of growing interest (Hudgens and Halloran,, 2008; Aronow and Samii,, 2017; Eckles et al.,, 2017; Athey et al.,, 2018). For many target effect estimands, interference should be minimized to obtain unbiased effect estimates. Here, if we have some a priori knowledge of how interference is likely to occur, we can use restricted randomization to rule out allocations that are likely to generate large amounts of interference.
More broadly, the core task in experimental design is to set up the study such that it generates a trustworthy answer to the study’s motivating causal question. We introduce Inspection-Guided Randomization (IGR), a flexible and transparent framework for restricted randomization in which the study analyst inspects the desirability of candidate randomization allocations, then constrains the candidate space to produce less biased and more precise estimates of the target estimand. The flexibility in IGR exists along two fronts. First, IGR permits targeting diverse design desiderata beyond those that relate to balance, thus expanding the scope of use cases for restricted randomization. Second, IGR incorporates a checkpoint that encourages the study analyst to carefully evaluate, and potentially adapt, the construction of the restricted assignment mechanism. IGR promotes transparency by enforcing that the set of acceptable treatment allocations be pre-registered in the trial protocol prior to running the study.
This work is organized as follows. In Section 2, we review prior work in restricted randomization and experimental design, then highlight the contributions we make through IGR. In Section 3, we present each step of the IGR framework in detail. In Section 4, we use illustrative vignettes with simulations to show how IGR can be applied to the design of group formation experiments and experiments with interference. In Section 5, we summarize the benefits of using IGR and propose directions for future work.
2 Related Work
Blocking and Matching
Blocking is a form of restricted randomization that rules out allocations that are imbalanced on one or more blocked covariates and has been used widely in the design of experiments (Fisher,, 1992; Box et al.,, 2005). Matched pair designs use matching algorithms to pair experimental units on the basis of a metric of covariate similarity, thus facilitating balance on multiple categorical and continuous covariates at once (Greevy et al.,, 2004; Xu and Kalbfleisch,, 2010; Imai et al.,, 2009). More recently, efficient blocking algorithms have been developed that extend matched-pair designs and allow for balanced blocks of flexible sizes (Higgins et al.,, 2016).
Tightening, Constraining, and Re-randomizing
An alternative approach to blocking and matching involves first generating a candidate set of allocations according to a pre-specified mechanism, then selecting one allocation from among this set as the official randomization to deploy. Tukey referred to this procedure as “tightening” the clinical trial (Tukey,, 1993), and Moulton later formalized a “covariate-constrained randomization” design for cluster-randomized trials (Moulton,, 2004). There has also been work in using constrained randomization in multi-arm settings with the use of multi-arm balance metrics (Ciolino et al.,, 2019; Watson et al.,, 2021; Zhou et al.,, 2021). With constrained designs, inference can be done via a randomization test over the candidate allocations.
Instead of listing all possible allocations at once, re-randomization repeatedly randomizes until producing the first allocation that meets pre-specified criteria (Rubin,, 2008; Cox,, 2009; Morgan and Rubin,, 2012). Re-randomization has been used beyond two-arm study designs in experiments with factorial designs (Branson et al.,, 2016) and in experiments on networks where units have correlated potential outcomes (Basse and Airoldi,, 2018). Although re-randomization can save computing power, it opens the door to p-hacking because the set of criteria-meeting allocations used to conduct a randomization test are sampled after study outcomes have been observed. Additionally, re-randomization assumes that the analyst’s first choice of restriction criteria produces a satisfactory design, when in reality the criteria may permit allocations that have poor properties for estimation and inference.
Optimal and Model-assisted Design
Kasy, (2016) argues that if the goal of the experiment is to minimize “disparity” between the estimated and true treatment effect (e.g., mean-squared error), one should deploy the single allocation that optimizes the expected loss over possible data-generating processes. Identifying an optimal allocation in this way, however, requires specifying a prior distribution for the loss function. Model-assisted design is similarly motivated, but optimizes the expected mean-squared error over random data samples from a specific data-generating process rather than for a distribution of data-generating processes (Basse and Airoldi,, 2018). Unlike with constrained randomization and re-randomization, when only a single optimal allocation is selected, randomization inference is no longer warranted.
Reproducible Science
Results of published studies are often not reproducible because they were generated through the use of questionable research practices aimed at “publication-worthiness” rather than scientific correctness and robustness (Baker,, 2016; Munafò et al.,, 2017; Ioannidis,, 2005). These practices include p-hacking, reworking research hypotheses after the results of a study have already been collected, and over-interpreting the data (Munafò et al.,, 2017; Ioannidis,, 2005). To promote reproducible and open science, there has been a push for pre-registering study protocols that hold researchers accountable to a planned study design and set of analyses (Munafò et al.,, 2017).
Our contributions
IGR unifies and builds upon ideas across several of the works above. The steps of IGR closely resemble those of covariate-constrained randomization (Moulton,, 2004) but, unlike covariate-constrained randomization, IGR recognizes that it can be beneficial to inspect allocations on the basis of multiple properties, particularly properties other than balance. IGR is also related to model-assisted design (Basse and Airoldi,, 2018), but whereas model-assisted design requires deriving a closed-form expression for mean-squared error, IGR can use domain-informed, approximate metrics when the expression for mean-squared error is unknown or difficult to derive. Furthermore, in line with efforts to make studies more reproducible, IGR includes a pre-registration step, which is not explicitly present in any existing restricted randomization framework. Finally, IGR places an emphasis on deliberately interrogating the quality of the restricted design and on making subsequent adaptations to the restriction as necessary. This adaptation step is either under-emphasized or absent from prior approaches to restricted randomization. In practice, we suspect that two common forms of adaptation will be: (i) fixing inspection metrics that do not correctly target the design desiderata (e.g., a balance metric that does not capture the kind of imbalance concerning to the study stakeholders) and (ii) adjusting the relative emphases placed on different design desiderata to better navigate trade-offs (e.g., in trials with interference, homophily can create tension between achieving covariate balance and minimizing interference, and which of the two desiderata is prioritized may need to be adjusted). We explore the latter form of adaptation in Vignette 2 (Section 4.2) and discuss further in Section 5.
3 Inspection-Guided Randomization
3.1 Setup and notation
We consider a setting with $N$ participants in a $k$-armed study, where the participants are indexed by $i=1,...,N$. Let $\mathbf{X}$ be the $N$-by-$p$ dimensional matrix representing the participants’ baseline covariates, where we denote the covariate vector for participant $i$ as $X_{i\cdot}$ and the vector of values for covariate $j$ as $X_{\cdot j}$. For a given participant $i$, we use the notation $Z_{i}$ to denote the participant’s treatment assignment, where $Z_{i}\in\{0,...,k-1\}$. Let the $N$-dimensional vector of treatments be denoted $\mathbf{Z}=(Z_{1},Z_{2},\dots,Z_{N})$. Since treatment is assigned at random, $Z_{i}$ is a random variable, and it follows that $\mathbf{Z}$ is a random vector. We denote a particular realization of a treatment assignment for individual $i$ as lowercase $z_{i}$ and likewise denote the intervention assignment vector lowercase $\mathbf{z}$. The probability that random vector $\mathbf{Z}$ takes on a particular value $\mathbf{z}$ is written $P(\mathbf{Z}=\mathbf{z})=p_{\mathbf{z}}$, and we refer to $P(\mathbf{Z})$ as the assignment mechanism. We call a particular realization of the intervention assignment vector a treatment allocation. Letting $\mathcal{Z}$ denote the collection of all $k^{N}$ possible values of $\mathbf{z}$, we have that $\sum_{\mathbf{z}\in\mathcal{Z}}p_{\mathbf{z}}=1$.
3.2 Designing experiments with IGR: A step-by-step framework
For clarity, we organize the IGR design framework into the following steps: 1. Specification, 2. Enumeration, 3. Restriction, 4. Evaluation & Adaptation, 5. Pre-registration, and 6. Randomization. We elaborate on each of these steps below.
1. Specification
The study analyst first works with the domain experts on the research team to specify the goals of the study in terms of the causal question of interest and the target effect estimand. Based on these goals, the team compiles desiderata for the study such that if the deployed randomization allocation were guaranteed to satisfy all desiderata, then the data collected would likely be conducive to credible causal inference.
The analyst then constructs a set of inspection metrics that are used to assess candidate allocations, where each metric measures the extent to which the allocation satisfies a particular desideratum (e.g., standardized mean difference as a metric to measure balance). We assume that these metrics are defined such that they map each allocation and observed covariates to a scalar value. We denote a collection of $L$ metrics as $m_{1},m_{2},\dots,m_{L}$ and denote $s_{1},s_{2},\dots s_{L}$ as the scalar outputs when these metrics are applied to inspect a particular allocation $\mathbf{z}$, where $s_{l}=m_{l}(\mathbf{z},\mathbf{X})\in\mathbb{R},\ l=1,2,\dots,L$.
The analyst also decides on an aggregator function, $\Theta$, that aggregates $s_{1},s_{2},\dots s_{L}$ into a single score that summarizes an allocation’s desirability. We write $\Theta(s_{1},s_{2},\dots,s_{L})=s$, where $s\in\mathbb{R}$. In certain settings, the desiderata may be conflicting, where scoring better for some desiderata necessarily entails worse scores on other desiderata. The aggregator can be used to navigate such desiderata trade-offs, for instance by incorporating a weighting scheme that prioritizes certain desiderata over others. Let $f$ be the function that maps an allocation $\mathbf{z}$ to its score by combining the outputs of each of the individual inspection metric via the aggregation function $\Theta$, that is $f(\mathbf{z},\mathbf{X})=\Theta(m_{1}(\mathbf{z},\mathbf{X}),m_{2}(\mathbf{z},\mathbf{X}),\dots,m_{L}(\mathbf{z},\mathbf{X})).$ We refer to $f$ as the fitness function.
Finally, the analyst specifies a restriction rule, $r$, that constrains the space of candidate treatment allocations on the basis of each allocation’s score, where $r(f(\mathbf{z},\mathbf{X}))=0$ means $\mathbf{z}$ is filtered out by setting $p_{\mathbf{z}}$ to zero and $r(f(\mathbf{z},\mathbf{X}))=1$ means $\mathbf{z}$ is deemed acceptable by setting $p_{\mathbf{z}}$ to a positive value between 0 and 1. For example, $r$ could be a thresholding rule, where only allocations with scores below some threshold $s^{*}$ are accepted.
2. Enumeration
The analyst enumerates a pool of unique candidate treatment allocations, $\mathcal{Z}_{pool}=\{\mathbf{z_{1}},\mathbf{z_{2}},\dots\mathbf{z_{M}}\}$. For a $k$-armed study, there are $k^{N}$ total possible allocations. When $N$ is large, it may be necessary to set $M<<k^{N}$. The enumeration of $\mathcal{Z}_{pool}$ can be done by sampling from any choice of assignment mechanism $P$, including Bernoulli randomization and complete randomization. In some cases, sampling from an assignment mechanism that is already restricted, such as a block randomization, can increase the efficiency of the enumeration step, since the mechanism inherently rules out some of the undesirable allocations. Similarly, it is also possible to use enumeration strategies that incorporate optimization routines and more efficiently search the sample space for allocations that have good fitness function scores (see Section 3.3).
3. Restriction
Applying the fitness function $f$ and restriction rule $r$ to the enumerated pool of candidate allocations yields the following assignment mechanism:
$P_{IGR}(\mathbf{Z}=\mathbf{z})=\begin{cases}0&\mathbf{z}\notin\mathcal{Z}_{pool}\\ 0&\mathbf{z}\in\mathcal{Z}_{pool},(r\circ f)(\mathbf{z},\mathbf{X})=0\\ p_{\mathbf{z}}&\mathbf{z}\in\mathcal{Z}_{pool},(r\circ f)(\mathbf{z},\mathbf{X})=1\end{cases}$
At this step, the assignment mechanism $P_{IGR}$ is proposed but not yet locked in.
4. Evaluation & Adaptation
Before proceeding and locking in the restricted assignment mechanism, the study analyst and domain experts pause to evaluate the fitness function $f$ and restriction rule $r$, making adaptations and improvements to them if necessary. Even if $f$ and $r$ were carefully identified in the Specification, they still may be flawed in unforeseen ways and could yield poor study designs. We highlight three dimensions along which $f$ and $r$ should be evaluated: discriminatory power, desiderata trade-offs, and over-restriction.
  1. 1.
    Discriminatory power: Fitness functions should provide meaningful differentiation between candidate allocations. If the fitness function used to score allocations produces a similar score for all candidates, then there would be no reason to expect that using a restricted assignment mechanism would confer any improvements in effect estimation. When confronted with a homogeneous set of scores, the fitness function may require redefinition, for instance through introducing additional inspection metrics.
  2. 2.
    Desiderata trade-offs: When there are multiple competing design desiderata, it may not be immediately clear whether and how to best construct the fitness function to weigh desiderata trade-offs. Visualizing the space of all candidate allocations by plotting the score outputs of different inspection metrics on separate axes can assist in understanding the nature of desiderata trade-offs. Then, visualizing the space of accepted allocations can assist in understanding whether the fitness function navigates such trade-offs as desired. If necessary, the analyst can make informed adjustments to the fitness function by modifying the aggregator function $\Theta$ that combines the inspection metric scores. In Section 4.2, we provide an example of how adjusting the prioritization weights for the individual inspection metrics can be used to navigate trade-offs between achieving covariate balance and minimizing interference between treatment arms.
  3. 3.
    Over-restriction: We consider two types of over-restriction. In the first type of over-restriction, the combination of fitness function and restriction rule may result in too few accepted allocations to obtain p-values with the desired level of granularity (see Section 3.4 for details on deriving p-values from randomization tests). When this occurs, the analyst may need to modify the restriction rule to be less aggressive or choose to apply a restriction rule that always selects the top $m$ allocations. Too few accepted allocations could also be a consequence of using an enumeration strategy that does not “hit” the correct regions of the space of all possible allocations.This issue can be addressed by increasing the size of the enumerated pool or by using stochastic optimization routines (see Section 3.3). In the second type of over-restriction, the set of accepted allocations may be too similar to one another, such that multiple units or groups of units are assigned to the same treatment arm across most accepted allocations. If we rely on an asymptotic p-value in testing the null hypothesis, then this kind of over-restriction can result in departures from the nominal Type I error rate (Moulton,, 2004). An exact p-value obtained through randomization inference circumvents this issue, but highly correlated allocations can also lead to increased mean-squared error and reduced power, regardless of whether randomization inference is used (Nordin and Schultzberg,, 2022; Krieger et al.,, 2020). Relaxing the restriction rule $r$ to permit additional allocations is one way to alleviate over-restriction issues.
5. Pre-registration
When the research team is satisfied with the choice of fitness function and restriction rule, they lock in the assignment mechanism $P_{IGR}$ and pre-register the fitness function $f$, the restriction rule $r$, and the pool of candidate allocations $\mathcal{Z}_{pool}$ in the trial protocol. If the number of candidate allocations is large, it may be more practical to pre-register the code, including the setting of a particular random seed, used to generate the pool of candidate allocations.
6. Randomization
Finally, the observed allocation $\mathbf{z}^{obs}\sim P_{IGR}$ is drawn and deployed.
3.3 IGR enhancements
3.3.1 Enhancement 1: Using genetic algorithms for efficient enumeration
Since $M<<k^{N}$, it is possible that many desirable, high-scoring allocations are overlooked and not enumerated as part of $\mathcal{Z}_{pool}$. Instead of using simple enumeration strategies, we may instead want to efficiently search through the space of all $k^{N}$ possible allocations to construct a more optimal $\mathcal{Z}_{pool}^{(*)}$. One approach to construct $\mathcal{Z}^{(*)}_{pool}$ is to use genetic algorithms (Mitchell,, 1998). The initial $\mathcal{Z}_{pool}$ is viewed as a population of treatment allocations. These allocations “mate” in pairs to produce a new generation of allocations, where novel allocations are introduced through “mutation” (single entries of an allocation are changed) and “crossover” (segments of two parent allocations are swapped). The individuals in this generation are scored using the specified fitness function $f$, and those with the best scores survive and breed to produce the next generation. We can repeat this process over $T$ generations to obtain $\mathcal{Z}_{pool}^{(*)}$. Henceforth, we use “IGRg” to refer to IGR enhanced by the use of genetic algorithms in the enumeration step.
3.3.2 Enhancement 2: Adding mirror allocations
Assuming no interference, the difference-in-means estimate is unbiased for the average treatment effect when the assignment mechanism satisfies the mirror property. For a binary intervention, the mirror property is satisfied if $P(Z_{i}=0)=P(Z_{i}=1)$ for all $i=1,\dots,N$. Suppose we enumerated all possible allocations in the Enumeration step. If we specify fitness function $f$ and restriction rule $r$ to be symmetric, such that $(r\circ f)(\mathbf{X},\mathbf{z}))=(r\circ f)(\mathbf{X},\mathbf{1}-\mathbf{z}))$, the mirror property will be satisfied by $P_{IGR}$. However, because we may not be able to enumerate all possible allocations, satisfaction of the mirror property is not guaranteed. Provided $f$ and $r$ are symmetric, we can update $P_{IGR}$ by ensuring that, for every allocation that is accepted, its mirror allocation is also accepted. Let $\tilde{\mathcal{Z}}_{pool}=\{1-\mathbf{z}\ |\mathbf{z}\in\mathcal{Z}_{pool}\}$ denote the set of mirror allocations for each allocation in $\mathcal{Z}_{pool}$. The assignment mechanism $\tilde{P}_{IGR}$ satisfies the mirror property, where
$\tilde{P}_{IGR}(\mathbf{Z}=\mathbf{z})=\begin{cases}0&\mathbf{z}\notin\mathcal{Z}_{pool}\cup\tilde{\mathcal{Z}}_{pool}\\ 0&\mathbf{z}\in\mathcal{Z}_{pool}\cup\tilde{\mathcal{Z}}_{pool},(r\circ f)(\mathbf{X},\mathbf{z})=0\\ p_{z}&\mathbf{z}\in\mathcal{Z}_{pool}\cup\tilde{\mathcal{Z}}_{pool},(r\circ f)(\mathbf{X},\mathbf{z})=1\end{cases}$
We can also naturally extend this mirroring to a multi-arm experiment in which one arm is designated the control arm and the remaining are treated.
In some experiments, the design desiderata may not be symmetric. For example, in an experiment with interference, the interference itself can create bias that is not resolved by satisfaction of the mirror property. Thus, choosing an “asymmetric” $f,r$ pair that does not satisfy the mirror property may sometimes help with reducing interference bias. However, asymmetry can also result in a violation of marginal independence between the treatment assignment variable and the potential outcomes. If $f,r$ is not perfectly symmetric, but the difference between $(r\circ f)(\mathbf{X},\mathbf{z})$ and $(r\circ f)(\mathbf{X},\mathbf{1}-\mathbf{z})$ is small for all $\mathbf{z}$, then the analyst may consider still incorporating the mirror allocations $\mathcal{\tilde{Z}}_{pool}$. We note that, while marginal independence may no longer hold under asymmetry, the treatment assignment variable will always be conditionally independent of the potential outcomes, where the conditioning set consists of the covariates used to calculate the fitness score. Thus, if not all mirrors are included, then a form of covariate adjustment can be performed during the analysis phase.
3.4 Analyzing IGR experiments: Randomization inference
We take a design-based approach to inference, meaning we view the potential outcomes as fixed conditional on the study sample, so that the only randomness in the study arises through the randomness in the treatment assignment mechanism. In IGR specifically, randomness arises in Step 6. Randomization, where the observed allocation is sampled from the restricted assignment mechanism, $P_{IGR}$. Randomization then forms the basis for inference, and we can perform a Fisher randomization test to evaluate null hypotheses. Specifically, we compute the value of the test statistic under the null hypothesis for different possible assignment vectors. With IGR, the empirical null distribution for the test statistic is derived over the set of assignment vectors $\mathbf{z}^{j}$ in $\mathcal{Z}_{pool}$ that have $p_{\mathbf{z}_{j}}>0$. We compare the observed value of the test statistic for $\mathbf{z}^{obs}$ to the empirically derived null distribution of the test statistic to obtain a p-value.
4 Applying IGR: Two Vignettes
In this section, we use vignettes and simulated experiments to illustrate how IGR can be used to achieve design desiderata in two distinct experimental settings:
  1. 1.
    Group formation experiments. In a group formation experiment, the challenge is to minimize differences in most baseline covariates, while creating large contrasts in a salient covariate of interest. Individuals are randomized into different groups, typically to construct groups of certain compositions that are believed to have different responses to the intervention (Basse et al.,, 2024). For example, a group may be the set of college freshmen assigned to the same dorm room, and composition may be the number of individuals in a room who were “high academic achievers” in high school (Sacerdote,, 2001). We consider group formation trials where composition is defined in terms of a categorical salient attribute (e.g., high or low prior academic achievement) and is count-based (e.g., number of individuals with high academic achievement). A well-designed experiment in this setting has groups balanced on all observed covariates other than the salient attribute itself.
  2. 2.
    Experiments with interference. Interference occurs when the treatment assigned to one observational unit affects the outcome of a different unit. With many interventions, interference is expected. For example, consider a behavioral intervention that teaches participants new information or skills. Participants in the treatment arm who receive the treatment may share what they have learned with their friends, some of whom may be assigned to the control arm. A common target estimand in experiments with interference is the total average treatment effect, or the difference in outcomes if everyone were to receive the treatment compared to if everyone were to receive the control. A well-designed experiment in this setting assigns treatments such that those in the treatment (control) arm experience a universe that is as close as possible to an all-treated (all-control) universe while also balancing covariates across arms.
In each vignette, we simulate an experiment designed using IGR and benchmark the results against alternative randomization designs (see Appendix A for the general simulation pipeline). Code is available at https://github.com/wangmagg/inspection-guided-randomization.
4.1 Vignette 1: IGR for group formation experiments
In this first vignette, we illustrate how to use IGR in the design of a group formation experiment where we are interested in evaluating how group composition influences the effect of a group-level intervention. Various kinds of inter-group comparisons may be of interest, including between treated and control groups with the same compositions (to target the treatment effect under a fixed composition) and between treated and treated groups with different compositions (to target the effect of the composition under a fixed treatment). In addition, we may also want to make difference-in-differences comparisons between the treatment effect under one composition and the treatment effect under a different composition.
The design goals in this setting are two-fold. First, we would like to sort individuals into groups such that certain pre-specified group compositions are realized. We focus in particular on the case where composition is characterized by the proportion of the group that possesses a particular salient attribute. Second, we would like to maintain balance on all covariates other than those used to define composition. A key point of this vignette is to highlight that in the Enumeration step of IGR, we can use different methods to sample allocations. Here, we sample from a blocking-based group formation randomization assignment mechanism.
We first discuss the choices we make in the Specification and Enumeration steps to meet these goals, then present results from a simulated group formation experiment.
4.1.1 Specification and Enumeration for a group formation experiment
We begin by establishing notation and formally clarifying the design goals in a group formation experiment. Suppose we want to form $N_{G}$ total groups. Let $g_{i}\in\{1,\dots N_{G}\}$ denote the group assignment for individual $i$ and $\mathbf{g}$ denote the $N$-vector of group assignments. Let $a$ be the salient attribute used to define group composition. Let $x_{ia}=1$ for individuals who have this attribute and $x_{ia}=0$ for individuals who do not have this attribute. We then define the composition of group $h$, denoted $\rho_{a}(h)$, as the proportion of the group that possess attribute $a$, where
$\rho_{a}(\mathbf{g},h,\mathbf{X})=\frac{\sum_{i=1}^{N}\mathbbm{1}\{x_{ia}=1\}\mathbbm{1}\{g_{i}=h\}}{\sum_{i=1}^{N}\mathbbm{1}\{g_{i}=h\}}$ (4.1.1)
Suppose that we would like to form groups that have $L$ unique compositions, $\mathcal{C}=\{c_{a}^{1},c_{a}^{2},\dots c_{a}^{L}\}$. Furthermore, we seek to form at least two groups with each composition, such that, for every composition, there can be at least one group with that composition assigned to treatment and one assigned to control.
Specification
We specify the target effect estimands as the average treatment effects under particular group compositions of interest. We denote the potential outcome for individual $i$ in group $h$ under treatment $z_{h}$ and composition $c_{a}^{l}\in\mathcal{C}$ as $Y_{ih}(z_{h},c_{a}^{l})$. The target estimands are $\tau_{l}=\frac{1}{N}\sum_{i=1}^{N}Y_{i}(1,c_{a}^{l})-Y_{i}(0,c_{a}^{l})\ \text{for all}\ c_{a}^{l}\in\mathcal{C}$. We specify one inspection metric that inspects allocations for balance and one for whether desired group compositions are realized.
Inspection Metric 1 (Balance): Standardized mean difference and Mahalanobis distance are commonly used balance metrics in two-arm studies. Here, we extend these metrics to the group formation experiment setting. Let $\mathbf{z}\in\{0,1\}^{N_{G}}$ denote the vector of treatment assignments across groups. Since there are both random treatment assignments $\mathbf{z}$ and random group assignments $\mathbf{g}$, we use fitness functions of the form $f(\mathbf{z},\mathbf{g},\mathbf{X})$.
We define $m^{(\texttt{SumMaxAbsSMD-G})}$ and $m^{(\texttt{MaxMahalanobis-G})}$ as
$m^{(\texttt{SumMaxAbsSMD-G})}(\mathbf{z},\mathbf{g},\mathbf{X})$ $=\sum_{j=1,j\neq a}^{p}\max_{(h_{1},h_{2})\in\mathcal{G}\times\mathcal{G}}\frac{\left|\bar{X}_{\cdot j}^{(h_{1})}-\bar{X}_{\cdot j}^{(h_{2})}\right|}{\sqrt{\left(s_{\cdot j}^{(h_{1})}\right)^{2}+\left(s_{\cdot j}^{(h_{2})}\right)^{2}}}$ (4.1.2)
$m^{(\texttt{MaxMahalanobis-G})}(\mathbf{z},\mathbf{g},\mathbf{X})$ $=\max_{(h_{1},h_{2})\in\mathcal{G}\times\mathcal{G}}\left(\bar{\mathbf{X}}_{\setminus a}^{(h_{1})}-\bar{\mathbf{X}}_{\setminus a}^{(h_{2})}\right)S_{\setminus a}^{-1}\left(\bar{\mathbf{X}}_{\setminus a}^{(h_{1})}-\bar{\mathbf{X}}_{\setminus a}^{(h_{2})}\right)^{T}$ (4.1.3)
where $\mathcal{G}=\{0,1,2,\cdots,N_{G}-1\}$ is the set of group indices, $\bar{X}_{\cdot j}^{(g)}$ and $s_{\cdot j}^{(h)}$ are the mean and standard deviation of covariate $j$ in group $h$, $\mathbf{\bar{X}}_{\setminus a}^{(h)}$ is the vector of covariate means in group $h$ excluding the salient attribute $a$, and $S_{\setminus a}$ is the sample covariance matrix excluding $a$.
Inspection Metric 2 (Composition Realization): To ensure allocations have the desired group compositions, we define the metric $m^{(\texttt{DesiredComps})}$ as
$m^{(\texttt{DesiredComps})}(\mathbf{z},\mathbf{g},\mathbf{X})=\begin{cases}1&\rho_{a}(\mathbf{g},h,\mathbf{X})\in\mathcal{C}\ \text{for all}\ h\in\mathcal{G},\\ &\mathbbm{1}\{\gamma^{0}(\mathbf{z},\mathbf{g},\mathbf{X},c_{a}^{l})>0\}\cdot\mathbbm{1}\{\gamma^{1}(\mathbf{z},\mathbf{g},\mathbf{X},c_{a}^{l})>0\}=1\\ &\ \ \ \ \text{for all}\ c_{a}^{l}\in\mathcal{C}\\ 0&\text{otherwise}\end{cases}$ (4.1.4)
where $\gamma^{0}(\mathbf{z},\mathbf{g},\mathbf{X},c_{a}^{l})=\sum_{h\in\mathcal{G}}\mathbbm{1}\{\rho_{a}(\mathbf{g},h,\mathbf{X})=c_{a}^{l}\}\cdot(1-\mathbf{z_{h}})$ is the number of control groups with composition $c_{a}^{l}$ and $\gamma^{1}(\mathbf{z},\mathbf{g},\mathbf{X},c_{a}^{l})=\sum_{h\in\mathcal{G}}\mathbbm{1}\{\rho_{a}(\mathbf{g},h,\mathbf{X})=c_{a}^{l}\}\cdot\mathbf{z_{h}}$ is the number of treated groups with composition $c_{a}^{l}$.
Combining these metrics, we define the fitness function
$\begin{split}f^{(\texttt{SumMaxAbsSMD-GD})}(\mathbf{z},\mathbf{g},\mathbf{X})&=\Theta(m^{(\texttt{SumMaxAbsSMD-G})}(\mathbf{z},\mathbf{g},\mathbf{X}),m^{(\texttt{DesiredComps})}(\mathbf{z},\mathbf{g},\mathbf{X}))\\ &=\begin{cases}m^{(\texttt{SumMaxAbsSMD-G})}(\mathbf{z},\mathbf{g},\mathbf{X}))&m^{(\texttt{DesiredComps})}(\mathbf{z},\mathbf{g},\mathbf{X})=1\\ \infty&m^{(\texttt{DesiredComps})}(\mathbf{z},\mathbf{g},\mathbf{X})=0\end{cases}\end{split}$ (4.1.5)
The fitness function $f^{(\texttt{MaxMahalanobis-GD})}$ is defined analogously. Lastly, we specify the restriction rule $r(s)=0,s\geq s^{*}$, $r(s)=1,s<s^{*}$, where $s^{*}$ is the $(100\cdot\frac{m}{M})^{\text{th}}$-percentile of scores. By applying this rule, $m$ out of the $M$ enumerated allocations are accepted.
Enumeration
For simplicity, we assume that we want to form groups of equal sizes such that there are exactly two groups with each composition, one of which is assigned treatment and one assigned control. We furthermore assume that the number of individuals that possess the salient attribute $a$ permits forming the desired groups. Then, to enumerate a candidate allocation, we first split the study population into two blocks, $\mathcal{B}_{1}=\{i:X_{ia}=1\}$ and $\mathcal{B}_{2}=\{i:X_{ia}=0\}$ We sample $c_{a}^{j}\cdot\frac{N}{N_{G}}$ individuals from $\mathcal{B}_{1}$ and $(1-c_{a}^{j})\cdot\frac{N}{N_{G}}$ from $\mathcal{B}_{2}$ uniformly at random, without replacement, to place into each pair of groups with the composition $c_{a}^{j}$. For every pair of groups with the same composition, we assign one group to treatment and one group to control. Under this group formation randomization assignment mechanism, $m^{(\texttt{DesiredComps})}(\mathbf{z},\mathbf{g},\mathbf{X})$ is guaranteed to equal 1. We refer to this assignment mechanism as group formation randomization (GFR). Note that, if we were to instead sample candidate allocations from an assignment mechanism that is agnostic to group composition, it would require enumerating an extremely large number of allocations to obtain any allocations that have the desired group compositions.
4.1.2 Simulated group formation experiment with perceived gender stereotyping as a contextual moderator
Data generation
Our group formation experiment simulation is inspired by stereotype threat interventions that have shown heterogeneous effects based on the gender composition of a group, where composition determines whether or not the individuals in the group perceive the presence of gender stereotyping (Grover et al.,, 2017; Carroll et al.,, 2023). The data generation process is modeled based on an unpublished educational intervention study that tested how different types of feedback on class assignments affect students’ grades. In the first two weeks of the class, the researchers collected information about students (age, gender, and major) and early course assignments (scores on a homework). After the second week of class, the researchers aimed to assign all $N$ students in the class to $k$ treatment arms (feedback conditions) with similar observed characteristics.
We simulate the study data set to encode theory-based covariance between student characteristics (age, gender, and college major), latent traits (ability and confidence), and course performance (homework score and final exam grade). The encoded covariance structure creates conditions under which it is difficult to achieve balanced assignment to multiple groups. We let the salient attribute $a$ be the gender covariate, where $x_{ia}=1$ if individual $i$ is a man, and we define group composition $\rho_{a}(g)$ as the proportion of individuals in group $g$ that are men. We designate exam grade as the outcome variable. To generate the observed outcomes, we apply an additive treatment effect, with the magnitude of the treatment effect varying across compositions.
We perform simulations with $M=100,000$ enumerated allocations and $m=500$ accepted allocations, which is in line with the number of enumerated and accepted allocations in prior work (Li et al.,, 2016; Watson et al.,, 2021; Nordin and Schultzberg,, 2022). We compare applying IGR using both a balance and a composition realization metric to applying IGR using a composition realization metric only. The latter design is functionally equivalent to sampling from the group formation randomization assignment mechanism described above, so we simply refer to it as “GFR” in the results. Further details on the data-generating process and the parameter settings used in the simulations can be found in Appendix B.
Results
Using the fitness function $f^{(\texttt{SumMaxAbsSMD-GD})}$ or $f^{(\texttt{MaxMahalanobis-GD})}$ (4.1.2) in conjunction with a thresholding restriction rule passes visual inspections for the “Discriminatory Power” and “Over-restriction” checks in the Evaluation & Adaptation step of IGR (Figure 4.1.1). By using $f^{(\texttt{SumMaxAbsSMD-GD})}$ or $f^{(\texttt{MaxMahalanobis-GD})}$ (4.1.5) as the fitness function, only allocations that have the desired compositions are accepted, and those with large imbalance on covariates are filtered out (Figure 4.1.2). In contrast, when individuals are randomized into groups of particular compositions using group formation randomization only, i.e. we apply the inspection metric $m^{(\texttt{DesiredComps})}$ as the fitness function, covariates are less balanced. Notably, including a balance metric helps improve balance on “Homework”, which is correlated with the salient attribute,“Gender”, and thus is highly imbalanced under GFR due to the differing gender compositions across groups. In Table 4.1.1, we show that, as expected, including a balance metric in IGR reduces the root mean-squared error (RMSE) of treatment effect estimates compared to using a composition metric alone.
Figure 4.1.1: [Vignette 1, $\{c_{a}^{1},c_{a}^{2},c_{a}^{3}\}=\{0.5,0.3,0.7\},\ N=120$], Visualizations for the Evaluation & Adaptation step: discriminatory power and over-restriction. (top) For both $m^{(\texttt{SumMaxAbsSMD-GD})}$ and $m^{(\texttt{MaxMahalanobis-GD})}$, there is substantive differentiation in score across candidate allocations, both when the allocations are enumerated without ($\mathcal{Z}_{pool}$) and with the addition of genetic algorithms ($\mathcal{Z}^{*}_{pool}$). $\mathcal{Z}^{*}_{pool}$ excludes the worst-scoring allocations in the right tail of $\mathcal{Z}_{pool}$. (bottom) IGR and IGRg only slightly increase pairwise assignment correlation relative to GFR.
Figure 4.1.2: [Vignette 1, $\{c_{a}^{1},c_{a}^{2},c_{a}^{3}\}=\{0.5,0.3,0.7\},\ N=120$] Standardized mean difference across all pairwise group comparisons for accepted allocations under Group Formation Randomization (GFR) and IGR/IGRg. Groups 0 and 1 are 50 % men, 2 and 3 are 30% men, and 4 and 5 are 70 % men. “Max” refers to the maximum SMD across all pairwise group comparisons. IGR improves balance on the non-salient covariates; using genetic algorithms (IGRg) does not noticeably enhance these improvements. Since “Homework” is correlated with “Gender”, it exhibits imbalance under GFR; this imbalance is reduced by including a balance metric in IGR. “Ability” and “Confidence” are latent covariates and are not included as direct inputs to the IGR balance metric. IGR improves balance on “Ability”, even though it is unobserved, because it is tightly correlated with “Homework.” To allow for a wider genetic algorithm search in IGRg, we permitted group sizes to range from 19 to 21 individuals and relaxed the composition realization constraint for allocations in child generations.
Group Comparison 0 vs 1 $(c_{a}^{1}=0.5,\text{ effect size }d=0.3)$ 2 vs 3 $(c_{a}^{2}=0.3,\text{ effect size }d=0.5)$ 4 vs 5 $(c_{a}^{3}=0.7,\text{ effect size d }=0.1)$ RMSE %$\%\text{RMSE}_{GFR}$ Rej. Rate RMSE %$\%\text{RMSE}_{GFR}$ Rej. Rate RMSE %$\%\text{RMSE}_{GFR}$ Rej. Rate Design Fitness GFR 0.39$\pm$0.02 Ref 0.16$\pm$0.02 0.40$\pm$0.02 Ref 0.35$\pm$0.04 0.39$\pm$0.02 Ref 0.05$\pm$0.00 IGR MaxMahalanobis-GD 58.49$\pm$2.57 0.37$\pm$0.04 57.07$\pm$3.68 0.78$\pm$0.02 55.91$\pm$4.44 0.08$\pm$0.01 SumMaxAbsSMD-GD 49.90$\pm$3.60 0.45$\pm$0.02 48.41$\pm$3.32 0.89$\pm$0.03 48.21$\pm$1.50 0.08$\pm$0.01 IGRg MaxMahalanobis-GD 56.82$\pm$2.97 0.38$\pm$0.04 55.52$\pm$4.41 0.80$\pm$0.04 56.39$\pm$3.69 0.08$\pm$0.01 SumMaxAbsSMD-GD 48.83$\pm$2.69 0.46$\pm$0.03 47.61$\pm$3.91 0.90$\pm$0.04 47.82$\pm$3.03 0.08$\pm$0.01
Table 4.1.1: [Vignette 1, $\{c_{a}^{1},c_{a}^{2},c_{a}^{3}\}=\{0.5,0.3,0.7\},\ N=120$] Root mean-squared error (RMSE) in treatment effect estimates and rejection rate of sharp null for IGR/IGRg compared to GFR in a group formation experiment. IGR/IGRg with both a balance and a composition realization metric reduces RMSE across group comparisons compared to GFR (equivalently, IGR with a composition metric alone). Means and standard deviations are computed over 5 data samples.
4.2 Vignette 2: IGR for experiments with interference
In the second vignette, we illustrate how to use IGR for an experiment where we expect interference, which occurs when one individual’s treatment affects a different individual’s outcome (Rosenbaum,, 2007). Interference is often present with behavioral interventions, where information can diffuse or spill over along social ties (Valente,, 2012; Cai et al.,, 2015; Kim et al.,, 2015). When interference is present, a common target treatment effect estimand is the total average treatment effect, or a comparison of the average outcome if everyone were to receive treatment versus if everyone were to receive control (Hudgens and Halloran,, 2008). Failing to properly account for interference can make estimates of the total treatment effect both biased and imprecise (Eckles et al.,, 2017). This in turn can result in misleading conclusions (e.g., false nulls) and misinformed recommendations for policy-making (e.g., failure to implement an effective intervention).
A common approach to controlling interference is to use cluster randomization, where randomization is done on the cluster level such that all individuals in the same cluster receive the same treatment (Hayes and Moulton,, 2017). Given enough separation between clusters, there is likely no spillover between different treatment arms, and comparing treated to control clusters can give an unbiased estimate of the total average treatment effect. However, it may sometimes be difficult to identify clusters that are sufficiently separated to avoid any inter-cluster interference. In other cases, it may be possible to identify a few large clusters that are well-separated, such as villages that are far apart geographically. However, randomizing only a few clusters leads to low power, especially if there is strong intra-cluster correlation on prognostic covariates. Often, for clusters that form organically, individuals within a cluster are more similar than those in different clusters, so intra-cluster correlation is expected.
Within the IGR design framework, we can be more flexible with the way that we control interference. Leveraging prior knowledge, which may not necessarily be complete or fully correct, we can use an interference-targeting fitness function to control the amount of anticipated interference in a more precise way than with cluster randomization. Even if we do not have perfect knowledge of the interference structure (e.g., which individuals interact to produce spillover), we can still develop “coarse” metrics that measure interference.
In this section, we discuss the choices we make in the Specification and Enumeration steps of IGR, then present simulations based on a sexual assault prevention intervention.
4.2.1 Specification and Enumeration for an experiment with interference
Specification
We define the target effect estimand to be the total average treatment effect, $\tau_{TATE}=\frac{1}{N}\sum_{i=1}^{N}Y_{i}(\mathbf{1})-Y_{i}(\mathbf{0})$ where $\mathbf{1}$ is the all-ones vector (everyone is treated) and $\mathbf{0}$ is the all-zeros vector (everyone is control). We specify one inspection metric for interference and one for balance.
Inspection Metric 1 (Interference): We present two example metrics for controlling interference. The first can be used even with coarse knowledge of the interference structure whereas the second requires access to the interference network.
For interventions that spread via person-to-person interaction, the likelihood of such spread or spillover may decrease with geographic distance between individuals. While we may not know precisely which individual(s) will interact with which other individual(s), we can try to allocate treatments in such a way that no two individuals assigned to treatment and control are too close to one another geographically. Closeness can be defined, for example, in terms of where they live, where they attend school, and where they work. This line of reasoning motivates an inspection metric that measures the minimum distance between any two individuals assigned to treatment and control, where a larger minimum distance is more desirable. To stay consistent with smaller scores being better, we use the reciprocal of the minimum distance. Formally, we define inspection metric $m^{(\texttt{InvMinEuclideanDist})}$ as
$m^{(\texttt{InvMinEuclideanDist})}=\left(\min_{(i,j)\in[N]\times[N],z_{i}\neq z_{j}}||\mathbf{l}_{i}-\mathbf{l}_{j}||_{2}^{2}\right)^{-1}$ (4.2.1)
where $z_{i},z_{j}\in\{0,1\}$ are the treatment assignments for units $i$ and $j$, respectively and $\mathbf{l}_{i},\mathbf{l}_{j}\in\mathbb{R}^{2}$ are coordinates for $i$ and $j$. While we use Euclidean distance here, other distance functions can also be used. It is also possible to expand upon a simple distance function by incorporating known information about ease of travel, such as the presence or absence of interconnecting roads.
In some cases, we may have more detailed knowledge of a network of interactions between individuals. For instance, we can ask participants to list the friends with whom they communicate regularly and build a network based on each participant’s list of contacts (Cai et al.,, 2015; Kim et al.,, 2015). In the network interference literature, it is common to assume that an individual is “exposed” through the treatments of their neighbors in the network via a neighborhood exposure model (Aronow and Samii,, 2017; Manski,, 2013).
Suppose we have a priori knowledge of the underlying network and seek to limit the number of control individuals exposed to treatment. This motivates an inspection metric that measures the number of control individuals that are exposed to treatment under a particular allocation. Let $\mathbf{A}$ be an $N\times N$ adjacency matrix such that $A_{ij}=1$ if $i$ and $j$ are joined by an edge in the network and $A_{ij}=0$ otherwise. Let $\mathcal{N}(i)$ denote the neighborhood of $i$ and let $\mathbf{z}_{\mathcal{N}(i)}$ denote the treatment assignments for the individuals in this neighborhood. Let $h(\cdot)$ be an exposure function where $h(\mathbf{z}_{\mathcal{N}(i)})=1$ if individual $i$ is exposed to the treatment and $h(\mathbf{z}_{\mathcal{N}(i)})=0$ otherwise. For example, for a “one-neighbor” exposure model, $h(\mathbf{z}_{\mathcal{N}(i)})=\mathbbm{1}\{\sum_{j\in\mathcal{N}(i)}z_{j}>0\}$. Using this notation, we define the inspection metric $m^{(\texttt{FracCtrlExposed})}$ as
$m^{(\texttt{FracCtrlExposed})}(\mathbf{z},\mathbf{A})=\frac{\sum_{i=1}^{N}\mathbbm{1}\{h(\mathbf{z}_{\mathcal{N}(i)})\}\cdot(1-z_{i})}{\sum_{i=1}^{N}(1-z_{i})}$ (4.2.2)
Inspection Metric 2 (Balance): We assess covariate balance using the metric $m^{(\texttt{MaxMahalanobis})}$, which is defined similarly to $m^{(\texttt{MaxMahalanobis-G})}$ (4.1.2) except there is no salient attribute that is excluded.
Combining the interference and balance metrics, we define the fitness function
$\begin{split}f^{(\texttt{F+M})}(\mathbf{z},\mathbf{X})&=\Theta(m^{(\texttt{FracControlExposed})}(\mathbf{z},\mathbf{X}),m^{(\texttt{MaxMahalanobis})}(\mathbf{z},\mathbf{X}))\\ &=w_{1}\cdot m^{(\texttt{FracControlExposed})}(\mathbf{z},\mathbf{X})+w_{2}\cdot m^{(\texttt{MaxMahalanobis})}\cdot(\mathbf{z},\mathbf{X})\end{split}$ (4.2.3)
where $w_{1}$ and $w_{2}$ are scalar weights that can be used to tune whether there is heavier emphasis on the balance desideratum or the interference desideratum. The fitness function $f^{(\texttt{I+M})}$ is defined analogously, replacing $m^{(\texttt{FracControlExposed})}$ with $m^{(\texttt{InvMinEuclideanDist})}$. Lastly, we use the thresholding restriction rule as defined in Vignette 1 (4.1).
Enumeration
We sample from a complete randomization assignment mechanism that assigns an equal number of units to the treatment and control arms.
4.2.2 Simulated experiment with interference inspired by a sexual assault prevention intervention
Data generation
Our simulated experiment with interference is based on a trial of a sexual assault prevention intervention for adolescent girls conducted across schools in urban settlements in Kenya. During the intervention, the girls learned and practiced verbal and physical skills for resisting sexual assault (Baiocchi et al.,, 2017; Sarnquist et al.,, 2023).
An option for the design of this trial was to randomize the intervention at the settlement level, which would have strongly mitigated interference. However, the settlements were known to have a wide range of baseline rates of sexual assault, meaning randomizing on the settlement level would have made it difficult to achieve balance. Hence, the trial was instead randomized on the school level, and, because the settlements were densely populated, it is likely that girls attending different schools interacted with one another. The skills learned by girls assigned to receive the intervention may have been passed on to girls assigned to the control arm, resulting in interference.
In our simulation, we generate synthetic schools in five Kenyan settlements: Kibera, Mukuru, Huruma, Korogocho, and Dandora. Baseline covariates are sampled using a nested hierarchical model so that individuals attending the same schools are more similar than individuals attending different schools, and individuals within the same settlement are more similar than individuals across different settlements. We use a statistical network model to generate the interference network, where the probability two individuals are connected by an edge is inversely proportional to the $\gamma^{th}$ power of the Euclidean distance between the schools. To generate the observed outcome, we apply an additive treatment effect and assume that an individual is exposed to treatment if they directly receive treatment or if a certain fraction of their network neighbors receive treatment. Further details on the data-generating process and the parameter settings used in the simulations can be found in Appendix C.
Results
In this vignette, as was the case in Vignette 1 (Section 4.1), there are design desiderata that are in tension with one another — minimizing interference and maximizing balance. With IGR, the study analyst can leverage visualizations similar to those we offer in Figure 4.2.1 as part of the Evaluation & Adaptation step to gain insights into the nature of these types of desiderata trade-offs, which will likely vary by study. Even in our simple simulation, varying the parameter settings for the data-generating process produces unique trade-offs between balance and interference desiderata. Specifically, when the probability that two individuals are connected decays more strongly with the distance between their respective schools, i.e. the interference network is sparser, the trade-off between balance and interference becomes less pronounced (Figure 4.2.1).
We show through visualizations that the “Discriminatory Power” and “Over-restriction” checks in the Evaluation & Adaptation step are satisfied when we use $f^{(\texttt{F+M})}$ (4.2.3) as fitness function in conjunction with a thresholding restriction rule (Figure 4.2.2, top and bottom rows). We furthermore show that modifying how the interference and balance metrics are aggregated can be an effective means of tuning desiderata prioritization. By adjusting the weights $w_{1}$ and $w_{2}$ assigned to $m^{(\texttt{FracControlExposed})}$ and $m^{(\texttt{MaxMahalanobis})}$, respectively, we can prioritize accepting allocations with better interference properties or with better balance properties (Figure 4.2.2, middle row); achieving both simultaneously is difficult due to the trade-offs depicted in Figure 4.2.1. Without access to the underlying interference network, we can use a Euclidean distance-based proxy metric, $m^{(\texttt{InvMinEuclideanDist})}$ to measure interference. While $f^{(\texttt{I+M})}$ filters out a subset of high-interference allocations, it is not as effective as $f^{(\texttt{F+M})}$ (Appendix Figure D.1). Finally, prioritizing interference reduction in IGR decreases bias, particularly in denser graphs, but also generally leads to variance inflation (Table 4.2.1).
Figure 4.2.1: [Vignette 2, $N=4000$] Visualization of balance and interference desiderata trade-offs in candidate allocations. Each bin in the heatmap represents a count of the number of candidate allocations with $m^{(\texttt{MaxMahalanobis})}$ and $m^{(\texttt{FracControlExposed})}$ values that fall into that bin. Darker colors indicate a higher density of allocations. As $\gamma$ increases, the probability two individuals are connected decays more strongly with increasing Euclidean distance between the individuals’ respective schools. A downward-sloping trend is suggestive of a trade-off between achieving balance (small values on the x-axis) and reducing interference (small values on the y-axis). We observe a stronger trade-off for denser graphs.
Figure 4.2.2: [Vignette 2, sparsity $\gamma=0.5$, $N=4000$] Visualizations for the Evaluation & Adaptation step when $f^{(\texttt{F+M})}$ is used as fitness function. (top) For all three weighting schemes, there is differentiation in $f(\mathbf{z})$ across candidate allocations, whether they are enumerated naively ($\mathcal{Z}_{pool}$) or with the use of genetic algorithms ($\mathcal{Z}^{*}_{pool}$). (middle) Each point represents a single accepted allocation. Complete randomization (CR) is plotted for reference. When prioritizing interference reduction, we accept more imbalanced allocations. When prioritizing balance, we are less effective in reducing interference. The FracExpo score distribution is bimodal because FracExpo is asymmetric and tends to be larger for mirror allocations. (bottom) IGR/IGRg increases pairwise assignment correlation slightly when interference is prioritized.
Design Fitness Bias %$\%\text{Bias}_{CR}$ Var %$\%\text{Var}_{CR}$ Rej. Rate CR -0.41$\pm$0.14 Ref 0.61$\pm$0.10 Ref 0.15$\pm$0.04 IGR 0.25*MaxMahalanobis + 0.75*FracExpo 87.88$\pm$0.77 140.61$\pm$26.80 0.13$\pm$0.04 0.50*MaxMahalanobis + 0.50*FracExpo 93.99$\pm$0.42 22.35$\pm$3.15 0.55$\pm$0.17 0.75*MaxMahalanobis + 0.25*FracExpo 99.14$\pm$0.67 4.49$\pm$0.20 1.00$\pm$0.00 0.25*MaxMahalanobis + 0.75*InvMinEuclidDist 99.65$\pm$0.83 33.24$\pm$21.05 0.49$\pm$0.37 0.50*MaxMahalanobis + 0.50*InvMinEuclidDist 100.95$\pm$0.27 15.00$\pm$8.20 0.72$\pm$0.27 0.75*MaxMahalanobis + 0.25*InvMinEuclidDist 101.66$\pm$0.29 7.83$\pm$2.30 0.91$\pm$0.08 IGRg 0.25*MaxMahalanobis + 0.75*FracExpo 88.19$\pm$1.07 113.05$\pm$26.97 0.13$\pm$0.06 0.50*MaxMahalanobis + 0.50*FracExpo 96.21$\pm$0.87 9.36$\pm$1.53 0.88$\pm$0.10 0.75*MaxMahalanobis + 0.25*FracExpo 99.80$\pm$0.82 2.62$\pm$0.21 1.00$\pm$0.00 0.25*MaxMahalanobis + 0.75*InvMinEuclidDist 99.50$\pm$1.09 30.26$\pm$21.27 0.56$\pm$0.37 0.50*MaxMahalanobis + 0.50*InvMinEuclidDist 100.90$\pm$0.32 7.05$\pm$4.61 0.89$\pm$0.14 0.75*MaxMahalanobis + 0.25*InvMinEuclidDist 101.83$\pm$0.34 2.58$\pm$0.71 1.00$\pm$0.00
Table 4.2.1: [Vignette 2, sparsity $\gamma=0.5$, $N=4000$] Bias, sample variance, and rejection rate of effect estimates under IGR compared to complete cluster randomization (CR). Applying greater weight to the interference-related inspection metric is more effective for bias reduction. Applying greater weight to the balance-related inspection metric is more effective for reducing variance. While using $m^{(\texttt{InvMinEuclideanDist})}$ does not decrease bias for $\gamma=0.5$, it does for $\gamma=0.75$ (Supplemental Table D.1).
5 Discussion
We introduced Inspection-Guided Randomization as a flexible framework for performing restricted randomization with diverse design desiderata and used illustrative examples to showcase its usefulness in different experimental settings. The core motivation for IGR, as for any restricted randomization, is to prevent the possibility of choosing a bad randomization allocation that leads to an effect estimate that deviates dramatically from the true value of the estimand. For obviously bad allocations, there is little reason not to discard them; this has been well-understood since the early days of the experimental design literature. In a conversation between Cochran and Fisher (recounted by Rubin and Cochran), Fisher is reported to have said that, confronted with a randomized allocation that “exhibited substantial imbalance on a prognostically important baseline covariate,” he would “of course” rerandomize if the experiment had not started yet (Rubin,, 2008).
Historically, the focus of restricted randomization has been on achieving covariate balance in two-arm experiments that compare a single treatment intervention to a control or status quo. However, some interventions, often ones that are behavioral, need to be evaluated through more complicated restricted designs. Researchers across several disciplines, including medicine, education, psychology, and computer science, have recognized this need and have independently applied principles of the IGR framework to designing experiments, albeit without the formal structure of IGR (Murnane et al.,, 2023; Cho et al.,, 2021, 2022; Kizilcec et al.,, 2022; Zahrt et al.,, 2023). This suggests that IGR is conceptually straightforward and intuitive, and it also points to how an explicit IGR framework can help to bring consistency to existing efforts to expand the scope of restricted randomization.
In this work, we showed how the IGR framework can be used to design experiments in two settings where existing restricted randomization frameworks are not immediately applicable. First, in Vignette 1 (Section 4.1), we applied IGR to the design of group formation experiments, where effect sizes may exhibit dependence on group composition. Recent work has highlighted that peer context should be taken into consideration as a driver of treatment effect heterogeneity for behavioral and psychological interventions (Yeager et al.,, 2019; Walton and Yeager,, 2020; Kizilcec et al.,, 2020; Kizilcec and Cohen,, 2017). Understanding how an intervention depends on peer context has important implications for understanding causal mechanisms, for devising policies that target certain contexts, and for designing concurrent interventions that can be used to manipulate contexts. IGR can also be broadly compatible with definitions of peer contexts beyond the one used in Vignette 1, such as those that involve more than one attribute or those that are continuous functions of peer characteristics.
In Vignette 2 (Section 4.2), we applied IGR to the design of experiments with interference. Experimental designs that seek to mitigate interference often do so through a form of cluster randomization, either over naturally existent clusters or over algorithmically detected ones (Ugander et al.,, 2013; Ugander and Yin,, 2023). Without IGR, the study analyst chooses between randomizing at the more aggregated level to limit interference, thus sacrificing power and covariate balance, or randomizing at the more granular level to increase efficiency, thus permitting more interference bias. With IGR, the analyst can construct fitness functions that combine both interference and balance metrics, thus reducing interference bias while maintaining efficiency. IGR can also be used even if we have limited knowledge of interference structure and need to rely on an approximate measure of interference. While we explored a specific setting where interference occurred via neighborhood exposure in a network of interactions, IGR can easily be applied to other interference structures by modifying the inspection metric used to measure the extent of anticipated interference.
Fundamentally, IGR compels the study analyst to invest time and care in the design phase of the experiment. The Specification and Evaluation & Adaptation steps in particular require considerable thoughtfulness, from specifying the study’s design desiderata, to developing the inspection metrics used to measure desiderata satisfaction, to constructing and evaluating the fitness function and restriction rule that characterize the restricted assignment mechanism. Attentiveness to design can produce large payoffs in the bias and precision of effect estimates. While using a tailored estimator in the analysis phase can sometimes produce similar payoffs, focusing efforts on the design phase has additional merits. For one, even the most clever estimators may not be able to salvage a study that was performed under an extremely poor choice of randomization allocation. In addition, as previously mentioned, the Pre-registration step precludes statistical hacking, as long as the analyst is held accountable to doing randomization inference over a pre-registered space of randomization allocations.
We call out the Evaluation & Adaptation step of IGR as an especially critical component of the framework. In some sense, the inclusion of the Evaluation & Adaptation step is a recognition that the study analyst is not all-knowing and hence may not select a good fitness function and restriction rule in their first attempt. In this work, we highlighted three issues to consider in Evaluation & Adaptation: desiderata trade-offs, discriminatory power, and over-restriction. For example, in Vignette 2, the analyst may initially choose to assign higher weight to the interference metric than to the balance metric in order to prioritize reducing interference bias (Figure 4.2.2). After observing the costs that this choice of prioritization incurs in balance, however, the analyst may choose to modify the weighting scheme to instead apply equal weights to both metrics. If the analyst notices that the chosen fitness function produces homogeneous scores or that the restricted assignment mechanism accepts allocations that are highly correlated, then they can make informed updates to the fitness function and to the restriction rule. These iterative refinements push the design towards one that is better and more likely to yield credible inference. As mentioned in Related Works (Section 2), an emphasis on detecting and correcting flawed restricted assignment mechanisms is missing from the standard re-randomization framework.
There are several interesting directions for future work on IGR. Since the fitness function in IGR is hand-crafted, it may be useful to develop theory on how similar the specified fitness function needs to be to the true conditional mean-squared error in order for IGR to produce better effect estimates than naive randomization strategies. When the fitness function is not a faithful proxy, then analysts may choose to use a non-IGR design. Alternatively, analysts could choose to improve the fitness function by using additional sources of data and domain knowledge. If experiments consist of multiple waves or are preceded by a pilot study, then there is opportunity to use information collected from the prior wave or from the pilot study to refine the fitness function used in the next wave or in the actual trial. In Vignette 2, for example, researchers could use a pilot study to estimate the interference network structure within a subsample of the study participants, then extrapolate this network to the entirety of the study sample. Fitness functions for new experiments could potentially also leverage data from prior historical experiments that evaluated different but comparable interventions in similar settings. Lastly, to make IGR as accessible as possible by study analysts, good software for implementation will be useful; this will likely include a dashboard interface that allows users to easily perform each of the steps of IGR, including interacting with visualizations.
Appendix A Simulation Pipeline
Simulation Pipeline
  1. 1.
    Generate baseline covariates according to a vignette-specific data-generating process.
  2. 2.
    “Design” the experiment, either using IGR or using a benchmark design strategy.
    1. (a)
      For IGR: Enumerate a pool of $M$ candidate allocations and apply a thresholding restriction rule where the threshold is defined as the $m^{th}$ smallest score. Specifically, keep the top $\frac{m}{k}$ scoring allocations and their corresponding $\frac{(k-1)\cdot m}{k}$ mirror allocations (regardless of score) as acceptable allocations for the experiment.
    2. (b)
      For benchmark design strategies: Generate $m$ possible allocations under the benchmark design. While this may not be inherently necessary for executing the design itself, doing so facilitates comparability of the design with IGR.
  3. 3.
    “Run” the experiment under each of the $m$ accepted allocations.
    1. (a)
      For each allocation, generate the observed outcomes. Then, estimate the target treatment effect estimand and use a randomization test to compute an exact p-value for a test of no effect.
    2. (b)
      Estimate bias, mean-squared error, and rejection rate using the treatment effect estimates and p-values across the $m$ re-runs.
Appendix B Vignette 1 Data-generating Process
We simulate the study data set to encode theory-based covariance between student characteristics (age, gender, and college major), latent traits (ability and confidence), and course performance (homework score and final exam grade). In particular, course performance is assumed to vary with students’ latent ability and confidence, which in turn depend on age, gender, and major. Ability is modeled to increase with age and is higher for some majors than other majors. In the context of STEM majors where the majority of students are men, we model confidence to be higher among men than women. We also model also higher confidence for certain STEM-related majors.
Let $\mathbf{x}_{i}=\begin{pmatrix}age_{i}&gender_{i}&major_{i}&hw_{i}&exam_{i}\end{pmatrix}^{T}$ denote observed baseline covariates for individual $i$ and $\mathbf{u}_{i}=\begin{pmatrix}ability_{i}&confidence_{i}\end{pmatrix}^{T}$ denote latent, unobserved covariates. To generate $gender_{i}$, we randomly set half of the study sample to be men. Although having a fixed gender composition in the study population is not strictly necessary, it makes it easier implement a group formation strategy that assembles groups with certain gender compositions. In a real-world setting, we could construct a study population with a fixed gender ratio by using an enrollment approach that intentionally targets this ratio. We use the following data-generating process for the remaining covariates:
$age_{i}$ $\sim\text{Unif}(19,25)$
$major_{i}$$\sim\{1,2,3\}\text{ with probability }\{0.5,0.3,0.2\}$
$ability_{i}$ $=(age_{i}-\bar{age})/\sigma_{age}-0.5\cdot\mathbbm{1}\{gender_{i}=1\}+\mathbbm{1}\{major_{i}=1\}+\varepsilon$
$confidence_{i}$ $=\mathbbm{1}\{gender_{i}=1\}+\mathbbm{1}\{major_{i}=2\}+\varepsilon$
$hw_{i}$ $=ability_{i}+\epsilon$
$exam_{i}$ $=ability_{i}+confidence_{i}+\varepsilon$
$\varepsilon$ $\sim N(0,1)$
We designate exam grade the outcome variable and generate the observed outcome $Y_{i}^{obs}$ for individual $i$ under a binary treatment $z_{i}\in\{0,1\}$ and composition index $c(i)$ as
$Y_{i}^{obs}=\text{exam}_{i}+\mathbf{e}_{c(i)}^{T}\boldsymbol{\tau}z_{i}$ (B.1)
where $\mathbf{e}_{c(i)}\in\mathbb{R}^{l}$ is a vector that has a 1 in the $c(i)^{th}$ entry and a 0 everywhere else, and $\boldsymbol{\tau}\in\mathbb{R}^{l}$ is a vector of treatment effects. Entry $j$ in $\boldsymbol{\tau}$ can be interpreted as the effect of the intervention for the composition $r_{a}^{j}$. We run simulations with groups of composition types $\{c_{a}^{1},c_{a}^{2},c_{a}^{3}\}=\{0.5,0.3,0.7\}$ (that is, 50% male, 30% male, and 70% male), with 20 individuals per group. In the Enumeration step of IGR, we enumerate a pool of size $M=100,000$ using a group formation randomization, then apply a thresholding decision rule to keep the $m=500$ allocations with the best scores.
Appendix C Vignette 2 Data-generating Process
We use GoogleEarth to determine an approximately central coordinate for each of the five settlements, then sample school coordinates in a grid around the settlement center. We generate covariates $\tilde{\mathbf{x}}_{i}=\begin{pmatrix}\tilde{x}_{i1},\tilde{x}_{i2}\end{pmatrix}^{T}$for each individual as follows:
$\begin{split}\tilde{x}_{ij}&\sim N(\mu_{s(i),j},\sigma_{individual,j})\\ \mu_{s(i),j}&\sim N(\mu_{t((s(i)),j},\sigma_{school,j})\\ \text{for }j&=1,2\end{split}$ (C.1)
where $s(i)$ denotes the school that $i$ attends and $t(s(i))\in\{1,2,3,4,5\}$ denotes the settlement in which $s(i)$ is located. We manually set $\mu_{t,j}$ to fixed values for each settlement $t\in\{1,2,3,4,5\}$ and each $j\in\{1,2\}$. Similarly, we also manually set $\sigma_{individual,j},\sigma_{school,j}$ for each $j\in\{1,2\}$. To introduce homophily, such that neighbors in the interference network are more similar in their covariates, we define $\mathbf{x}_{i}$ to be the sum of $\tilde{x}_{i}$ and $\tilde{x}_{j}$ for all $j$ that are $i$’s neighbors in the network. Formally,
$\mathbf{x}_{i}=\mathbf{A}\tilde{\mathbf{X}}$ (C.2)
where $\mathbf{A}$ is an adjacency matrix for the interference network and $\mathbf{\tilde{X}}=\begin{pmatrix}\tilde{x}_{1}&\tilde{x}_{2}&\cdots&\tilde{x}_{n}\end{pmatrix}^{T}$. We sample edges for the underlying interference network using the following model:
$P(A_{ij}=1)$ =$=\begin{cases}\theta^{w}&s(i)=s(j)\\ \theta^{b1}\cdot\tilde{d}(i,j)^{-\gamma}&s(i)\neq s(j),t(i)=t(j)\\ \theta^{b2}\cdot\tilde{d}(i,j)^{-\gamma}&s(i)\neq s(j),t(i)\neq t(j)\\ \end{cases}$ (C.3)
$\tilde{d}(i,j)$ $=\frac{||\mathbf{l}_{s(i)}-\mathbf{l}_{s(j)}||_{2}^{2}}{\min_{(i,j)\in[N]\times[N],s(i)\neq s(j)}||\mathbf{l}_{s(i)}-\mathbf{l}_{s(j)}||_{2}^{2}}$ (C.4)
where $A_{ij}$ is the $(i,j)$-th entry of the adjacency matrix, $s(\cdot)$ denotes school membership, $t(s(\cdot))$ denotes settlement membership, and $\mathbf{l}_{s(\cdot)}$ denotes the school’s 2-dimensional latitude and longitude coordinates. $\theta^{w}$, $\theta^{b1}$, and $\theta^{b2}$ are scalar factors. We distinguish between $\theta^{b1}$ and $\theta^{b2}$ because individuals who attend different schools within the same settlement may be more likely to be connected than individuals who attend different schools in different settlements, even if the schools are the same distance apart.
To generate the observed outcome $Y_{i}^{obs}$ for individual $i$ under global treatment assignment vector $\mathbf{z}^{obs}\in\{0,1\}^{N}$, we use the following model, with $\tau$ capturing the direct effect of an individual’s own treatment and $\delta$ capturing the additive spillover effect:
$Y_{i}^{obs}=\mathbf{X}\boldsymbol{\beta}+\tau z_{i}+\delta\mathbbm{1}\left\{\sum_{j\in\mathcal{N}(i)}z_{j}>q\cdot|\mathcal{N}(i)|\right\}(1-z_{i})$ (C.5)
Note that, under this model, the total average treatment effect $\tau_{TATE}$ is exactly equal to $\tau$. Furthermore, if we use an experimental design that assigns a fixed number of units $N_{c}$ to control, then the bias of a difference-in-means estimator is
$E[\hat{\tau}-\tau]=-\delta\cdot\frac{1}{N_{c}}\sum_{i=1}^{N}\mathbb{E}\left[\mathbbm{1}\left\{\sum_{j\in\mathcal{N}(i)}z_{j}>q\cdot|\mathcal{N}(i)|\right\}(1-z_{i})\right]$
where the expectation is taken over all possible allocations. Thus, in order to reduce interference bias, we should rule out allocations with many exposed control units.
We run simulations for ($\boldsymbol{\mu}_{\cdot 1},\boldsymbol{\mu}_{\cdot 2})=\left(\begin{pmatrix}0.25&0&0.05&0&1\end{pmatrix},\begin{pmatrix}0.25&0.75&0&0.25&0\end{pmatrix}\right)$;
$(\sigma_{individual,j},\sigma_{school,j})=(0.1\cdot sd(\mu_{\cdot j}),0.2\cdot sd(\mu_{\cdot j}))$; $\gamma\in\{0.25,0.50,0.75\}$; ($\theta^{w},\theta^{b1},\theta^{b2})=(0.2,0.1,0.01)$; $\boldsymbol{\beta}=\begin{pmatrix}1&1\end{pmatrix}^{T}$; $q=0.25$; and $\tau=\delta=0.3\cdot SD(\mathbf{X}\boldsymbol{\beta})$. For IGR, we enumerate the pool of candidate allocations using a complete cluster randomization design, then apply a thresholding decision rule. In the main text, we show results for an enumerated pool of size $M=100,000$ and an accepted set of size $m=500$. Results for $M\in\{5000,1000,25000,50000,1000000,200000\}$ and $m\in\{100,500,1000\}$ are in Section D.
Appendix D Additional Results from Vignette 2
Figure D.1: [Vignette 2, sparsity $\gamma=0.5$, $N=4000$] Visualizations for the Evaluation & Adaptation step when $f^{(\texttt{I+M})}$ is used as the fitness function. (top row) $m^{(\texttt{InvMinEuclideanDist})}$ is only takes on a handful of distinct values, so discriminatory power is driven by $m^{(\texttt{MaxMahalanobis})}$. (middle row) While $m^{(\texttt{InMinEuclideanDist})}$ is coarser than $m^{(\texttt{FracCtrlExposed})}$, it still helps rule out the worst allocations. (bottom row) Pairwise assignment correlation under IGR and IGRg is comparable to that of CR.
Bias %$\%\text{Bias}_{CR}$ Var %$\%\text{Var}_{CR}$ Rej. Rate Design Fitness CR -0.12$\pm$0.05 Ref 0.48$\pm$0.07 Ref 0.24$\pm$0.03 IGR 0.25*MaxMahalanobis + 0.75*FracExpo 85.53$\pm$1.85 82.33$\pm$23.82 0.32$\pm$0.15 0.50*MaxMahalanobis + 0.50*FracExpo 90.62$\pm$2.42 17.01$\pm$2.85 0.89$\pm$0.06 0.75*MaxMahalanobis + 0.25*FracExpo 96.00$\pm$2.06 4.90$\pm$0.25 1.00$\pm$0.00 0.25*MaxMahalanobis + 0.75*InvMinEuclidDist 95.68$\pm$1.24 32.96$\pm$19.87 0.66$\pm$0.30 0.50*MaxMahalanobis + 0.50*InvMinEuclidDist 97.45$\pm$1.43 14.96$\pm$7.74 0.90$\pm$0.12 0.75*MaxMahalanobis + 0.25*InvMinEuclidDist 98.68$\pm$1.73 7.83$\pm$2.32 1.00$\pm$0.01 IGRg 0.25*MaxMahalanobis + 0.75*FracExpo 84.63$\pm$2.70 62.68$\pm$10.89 0.35$\pm$0.07 0.50*MaxMahalanobis + 0.50*FracExpo 90.46$\pm$3.16 9.70$\pm$1.17 0.98$\pm$0.01 0.75*MaxMahalanobis + 0.25*FracExpo 96.41$\pm$2.25 2.09$\pm$0.27 1.00$\pm$0.00 0.25*MaxMahalanobis + 0.75*InvMinEuclidDist 94.62$\pm$0.92 25.24$\pm$17.51 0.74$\pm$0.24 0.50*MaxMahalanobis + 0.50*InvMinEuclidDist 97.40$\pm$1.65 6.84$\pm$4.40 0.98$\pm$0.03 0.75*MaxMahalanobis + 0.25*InvMinEuclidDist 98.87$\pm$1.17 2.75$\pm$0.73 1.00$\pm$0.00
Table D.1: [Vignette 2, sparsity $\gamma=0.75$, $N=4000$] Bias, sample variance, and rejection rate of effect estimates under IGR compared to complete cluster randomization (CR). Applying greater weight to the interference-related inspection metric is more effective for bias reduction. Applying greater weight to the balance-related inspection metric is more effective for reducing variance. The coarser interference metric $m^{(\texttt{InvMinEuclideanDist})}$ helps to decrease interference bias, albeit not as effectively $m^{(\texttt{FracExpo})}$.
Figure D.2: [Vignette 2, sparsity $\gamma=0.5$, $N=4000$, $w_{balance}=0.5$, $w_{interference}=0.5$] Bias, variance, and rejection rate versus number of enumerated candidate allocations $M$ and the number of accepted allocations $m$. Thin lines represent results from a single data sample. Thick lines represent means taken across five data samples. When $m^{(\texttt{FracExpo})}$ is used as the interference metric, increasing $M$ results in greater bias and variance reductions relative to complete randomization (CR). When $m^{(\texttt{InvMinEuclidDist})}$ is used as the interference metric, increasing $M$ results in a slightly greater bias reduction and a more noticeable gain in variance reduction. Rejection rate increases as $M$ increases for both interference metrics.
Figure D.3: [Vignette 2, sparsity $\gamma=0.5$, $N=4000$] Bias, variance, and rejection rate versus magnitude of the weight applied to the balance metric, $m^{(\texttt{MaxMahalanobis})}$. Across plotted points, the weight applied to the interference metric is equal to 1 - $w_{\text{MaxMahalanobis}}$. As $w_{\text{MaxMahalanobis}}$ increases, IGR yields less bias reduction compared to complete randomization (CR). IGR can increase bias compared to CR when $w_{\text{MaxMahalanobis}}$ is large ($\gtrapprox 0.8$ for $m^{(\texttt{FracExpo})}$ and $\gtrapprox 0.4$ for $m^{(\texttt{InvMinEuclidDist})}$). As $w_{\text{MaxMahalanobis}}$ increases, IGR yields greater variance reduction compared to CR. IGR can increase variance compared to CR when $w_{\text{MaxMahalanobis}}$ is small ($\lessapprox 3$ for $m^{(\texttt{FracExpo})}$). Rejection rate increases as $w_{\text{MaxMahalanobis}}$ increases.
Appendix E Vignette 0: Multi-arm Experiment
In this additional vignette, we illustrate how to use IGR in the design of a multi-arm experiment, in a manner similar to Ciolino et al., (2019), Watson et al., (2021), and Zhou et al., (2011). In a multi-arm experiment, we simultaneously evaluate multiple interventions under a single trial protocol. To estimate treatment effects, we make pairwise comparisons between arms, often between each intervention arm and a single control arm. The design desideratum is to allocate treatments such that covariates are balanced across arms. We first discuss the choices we make in the Specification and Enumeration steps of IGR, then present results from a simulated experiment of a multi-arm classroom-based intervention.
E.1 Specification and Enumeration for designing a multi-arm experiment
Specification
We assume there is one control arm, and the remaining arms are treated with different interventions. We define the target effect estimands to be the average treatment effects comparing each of the treatments to the control, $\tau_{h}=\frac{1}{N}\sum_{i=1}^{N}Y_{i}(h)-Y_{i}(0)$ for all $h=1,\dots k-1$, where we assume arm 00 is the control arm and arms $11,\dots k-$ 1 are treated.
We assess covariate balance using the metrics $m^{(\texttt{MaxMahalanobis})}$ and $m^{(\texttt{SumMaxAbsSMD})}$, which are defined similarly to $m^{(\texttt{MaxMahalanobis-G})}$ and $m^{(\texttt{SumMaxAbsSMD-G})}$ (4.1.2) except there is no salient attribute that is excluded. Since there is only a single desideratum in this vignette, the aggregator function $\Theta$ is simply the identity function, and the fitness function is the inspection metric itself. Lastly, we use the thresholding restriction rule as defined in Vignette 1 (4.1).
Enumeration
To enumerate candidate allocations, we sample from a complete randomization assignment mechanism that assigns an equal number of individuals to each arm.
E.2 Simulated multi-arm experiment of educational feedback types
Data Generation
We use a similar data-generating process as the one detailed in Section B to generate baseline covariates, where all covariates are drawn in the same manner except for gender. In this vignette, we sample gender as $male_{i}\sim\text{Bern}(0.7)$ instead of setting a fixed gender composition in the study population. We again designate exam grade the outcome variable and define an outcome-generation process for the observed outcome $Y_{i}^{obs}$ for individual $i$ under treatment $z_{i}$ as
$Y_{i}^{obs}=\text{exam}_{i}+\sum_{j=1}^{k}\tau_{j}\mathbbm{1}\{z_{i}=j\}$ (E.1)
where $\{\tau_{j}\}_{j=1}^{k}\in\mathbb{R}^{k}$ are the treatment effects. The different arms correspond to different types of feedback on class assignments.
We run simulations of 4-arm and 6-arm experiments with 20 participants per arm. In the 4-arm experiment, one arm serves as control, and the remaining three are treated with interventions that have effect sizes of 0, 0.3, and 0.6. In the 6-arm experiment, one arm serves as control, and the remaining five are treated with interventions that have effect sizes of 0, 0, 0.1, 0.3, and 0.6. We benchmark IGR against complete randomization and blocked randomization with blocks derived via threshold blocking, a computationally efficient algorithm for forming near-optimal covariate-balanced blocks (Higgins et al.,, 2016).
Results
First, we show through visual assessments that the “Discriminatory Power” and “Over-restriction” checks in the Evaluation & Adaptation step are satisfied when we run IGR with $m^{(\texttt{SumMaxAbsSMD})}$ or $m^{(\texttt{MaxMahalanobis})}$ as the inspection metric (equivalently, fitness function) and a thresholding rule as the restriction rule. Specifically, to pass the “Discriminatory Power” check, we show that the scores for the enumerated candidate allocations are heterogeneous enough to make restriction on the basis of such scores reasonable. To pass the “Over-restriction” check, we show that pairs of individuals are no more likely to be assigned to the same treatment arm under IGR as they are under complete randomization. The same result holds when genetic algorithms are used in the enumeration step of IGR (IGRg).
Compared to complete randomization and threshold blocking, IGR leads to better balance on observed covariates, both globally and in each pairwise comparison between arms (Figure E.2). In Tables E.1 and E.2, we show that IGR reduces the root mean-squared error (RMSE) of treatment effect estimates. There is less RMSE reduction in the 6-arm experiment because achieving covariate balance becomes more difficult as the number of arms increases, so finding balanced allocations requires enumerating more candidate allocations.
Figure E.1: [Vignette 0, $k=4\text{ arms},N=80$] Visualizations for the Evaluation & Adaptation step: discriminatory power and over-restriction (top row) For both $m^{(\texttt{MaxMahalanobis})}$ and $m^{(\texttt{SumMaxAbsSMD})}$, there is substantive differentiation in score across candidate allocations, whether the allocations are enumerated naively ($\mathcal{Z}_{pool}$) or with the use of genetic algorithms ($\mathcal{Z}^{*}_{pool}$). The cutoff score (IGR s for $\mathcal{Z}_{pool}$ and IGRg s for $\mathcal{Z}^{*}_{pool}$) was used to filter out all but the 500 allocations with the smallest scores. While $\mathcal{Z}^{*}_{pool}$ excludes the worst-scoring allocations in the right tail of $\mathcal{Z}_{pool}$, it only somewhat improves the scores of the 500 accepted allocations. (bottom row) IGR and IGRg have similar pairwise assignment correlations compared to complete randomization (CR), whereas threshold blocking (QB) changes the pairwise assignment correlations.
Figure E.2: [Vignette 0, $k=4\text{ arms},N=80$] Standardized mean difference across all pairwise arm comparisons for the allocations accepted under Complete Randomization (CR), Threshold Blocking (QB), and Inspection-Guided Randomization (IGR, IGRg) (bottom) Standardized mean difference for each pairwise arm comparison. “Max” refers to the maximum SMD across all pairwise group comparisons. CR accepts some allocations with very large maximum standardized mean difference. QB accepts allocations that have smaller maximum standardized mean difference than CR. Using IGR with $m^{(\texttt{MaxMahalanobis})}$ and $m^{(\texttt{SumMaxAbsSMD})}$ as fitness functions rules out allocations with large maximum standardized mean difference across all covariates and is generally more effective than QB. Enumerating candidate allocations with genetic algorithms (IGRg) does not enhance balance compared to naive enumeration. “Ability” and “Confidence” are latent covariates and are not included as direct inputs to the IGR balance inspection metric. IGR nevertheless improves balance for these unobserved covariates because they are functions of the observed covariates.
Arm 1 (effect size d = 0) Arm 2 (effect size d = 0.3) Arm 3 (effect size d = 0.6) RMSE %$\%\text{RMSE}_{CR}$ Rej. Rate RMSE %$\%\text{RMSE}_{CR}$ Rej. Rate RMSE %$\%\text{RMSE}_{CR}$ Rej. Rate Design Fitness CR 0.39±0.04 Ref 0.04±0.01 0.39±0.03 Ref 0.14±0.01 0.39±0.05 Ref 0.43±0.02 QB 56.35±5.60 0.03±0.00 57.24±5.65 0.30±0.03 56.21±6.40 0.89±0.08 IGR MaxMahalanobis 49.07±1.49 0.02±0.01 48.39±3.69 0.38±0.03 48.77±2.74 0.95±0.02 SumMaxAbsSMD 45.79±3.45 0.02±0.00 44.31±2.81 0.43±0.05 45.45±3.35 0.98±0.01 IGRg MaxMahalanobis 46.75±2.49 0.02±0.00 46.23±3.39 0.39±0.03 46.28±2.21 0.97±0.01 SumMaxAbsSMD 43.26±3.47 0.02±0.00 42.77±2.37 0.43±0.05 42.92±5.59 0.98±0.01
Table E.1: [Vignette 0, $k=4\text{ arms},N=80$] Root mean-squared error (RMSE) in treatment effect estimates and rejection rate of sharp null for IGR/IGRg compared to complete randomization (CR) and threshold blocking (QB) in a multi-arm experiment. [$k=4$ arms, $N=80$] IGR and IGRg reduce RMSE across arms and increase the rejection rate in arms with non-zero treatment effects. Using IGR and IGRg confers more gains than using QB. Note that, to allow for a wider genetic algorithm search in IGRg, we permitted arm sizes to range from 19 to 21 individuals for allocations in child generations. Means and standard deviations are computed over 5 data samples.
RMSE %$\%\text{RMSE}_{CR}$ Rej. Rate Arm Design Fitness Arm 1 (effect size d = 0) CR 0.40±0.02 Ref 0.04±0.00 QB 60.84±4.71 0.03±0.01 IGR MaxMahalanobis 60.73±4.54 0.03±0.00 SumMaxAbsSMD 54.85±4.21 0.03±0.00 IGRg MaxMahalanobis 58.17±6.75 0.03±0.01 SumMaxAbsSMD 54.30±3.41 0.03±0.01 Arm 2 (effect size d = 0.15) CR 0.40±0.02 Ref 0.07±0.00 QB 60.08±3.08 0.09±0.01 IGR MaxMahalanobis 58.80±3.49 0.09±0.01 SumMaxAbsSMD 52.15±3.97 0.10±0.02 IGRg MaxMahalanobis 56.02±5.51 0.09±0.02 SumMaxAbsSMD 52.02±3.51 0.09±0.02 Arm 3 (effect size d = 0.3) CR 0.40±0.02 Ref 0.13±0.02 QB 60.91±3.62 0.26±0.03 IGR MaxMahalanobis 61.57±2.82 0.28±0.04 SumMaxAbsSMD 54.75±2.69 0.35±0.03 IGRg MaxMahalanobis 58.70±2.71 0.31±0.02 SumMaxAbsSMD 53.10±3.50 0.37±0.01 Arm 4 (effect size d = 0.45) CR 0.41±0.03 Ref 0.23±0.02 QB 58.95±5.57 0.58±0.04 IGR MaxMahalanobis 60.46±4.72 0.56±0.04 SumMaxAbsSMD 53.70±4.19 0.66±0.06 IGRg MaxMahalanobis 56.60±6.48 0.60±0.06 SumMaxAbsSMD 52.86±5.52 0.67±0.10 Arm 5 (effect size d = 0.6) CR 0.41±0.03 Ref 0.43±0.03 QB 59.02±0.78 0.86±0.04 IGR MaxMahalanobis 60.74±2.53 0.84±0.03 SumMaxAbsSMD 54.07±2.68 0.90±0.04 IGRg MaxMahalanobis 57.51±3.03 0.87±0.05 SumMaxAbsSMD 53.24±3.57 0.91±0.04
Table E.2: [Vignette 0, $k=6$ arms, $N=120$] Root mean-squared error (RMSE) in treatment effect estimates and rejection rate of sharp null for IGR/IGRg compared to complete randomization (CR) and threshold blocking (QB). Using IGR - SumMaxAbsSMD and IGRg - SumMaxAbsSMD results in lower RMSE than CR and QB. IGR - MaxMahalanobis and IGRg - MaxMahalanobis results in lower RMSE than CR, but not QB. Note that, to allow for a wider genetic algorithm search in IGRg, we permitted arm sizes to range from 19 to 21 individuals for allocations in child generations. Means and standard deviations are computed over 5 data samples.